# Elastic tensor-derived properties of composition-dependent disordered refractory binary alloys using DFPT

### v3

@@ 3.2 Rigid-ion elastic constants
Despite the lattice distortions in the binary alloys, the magnitudes of the relaxations lie within the range of the standard deviations and within 2% of the rigid-ion elastic constant values. These are also well within differences in experimental values, which can be as high as 10% [17, 18], depending on the experimental conditions and technique employed. Hence, the computationally faster rigid-ion elastic tensor calculations were performed for the additional compositions of A = 12.5 at.%, 37.5 at.%, 62.5 at.and 87.5 at.for all of the AB binary alloys. This allowed us to capture the effects of alloying on the elastic constants of BCC refractory binary alloys for an extended set of compositions.
@@
Figure 1: Composition-dependent averaged rigid-ion elastic constants $\bar{C}_{11}^{rigid}$, $\bar{C}_{12}^{rigid}$ and $\bar{C}_{44}^{rigid}$ (solid lines, Redlich-Kister polynomial fit to computed values (filled diamonds)), compared with experimental elastic constants (stars) for (a) MoNb [80]; (b) MoTa [81]; (c) MoW; (d) NbTa [82]; (e) WNb [83]; and (f) WTa [84].
@@
Fig. 1 compares the calculated rigid-ion elastic constants ($\bar{C}_{11}^{rigid}$, $\bar{C}_{12}^{rigid}$ and $\bar{C}_{44}^{rigid}$) for all compositions (filled diamonds) with experimental values (stars) for the binary alloys. The rigid-ion elastic constants are fitted to the Redlich-Kister polynomial given in Eq. (2.3) (solid line). For the same-group binary alloys (Figs. 1(c) and (d)), elastic constants are observed to vary linearly, similar to the empirical rule of Vegard’s law [85], while nonlinear behavior is observed for the different-group binary alloys. For the different-group binary alloys, the Ta-containing binary alloys (Figs. 1(b) and (f)) exhibit a greater degree of nonlinearity compared to the Nb-containing binary alloys (Figs. 1(a) and 1(e)). For $\bar{C}_{11}^{rigid}$ and $\bar{C}_{44}^{rigid}$ in the different-group binary alloys, a negative deviation from Vegard’s law is observed for the compositions with ${\rm X}>50$ at.where X is the element with lower elastic modulus value of the constituent elements. However, for $\bar{C}_{12}^{rigid}$, the trend is reversed, i.e., a positive deviation from Vegard’s law is observed for the compositions when ${\rm X}>50$ at.where X is the element with lower elastic modulus value of the constituent elements.
@@
Figure 2: Composition-dependent elastic moduli $B$, $G$ and $E$ computed using the Voigt method (Eqs. (11)–(13)), as a function of composition, compared with experimental $E$ (circles) for (a) MoNb [86, 87, 88]; (b) MoTa [86, 87, 88, 89]; (c) MoW [86, 87, 88]; (d) NbTa [86, 87, 88, 89]; (e) WNb [86, 87, 88]; and (f) WTa [86, 87, 88, 89]. Error bars correspond to the standard deviation about averaged experimental values.
@@ 3.2.1 Comparison with experiment
Experimental values at room temperature for all elastic constants were obtained from the literature as measured using high-frequency ultrasonic methods. For MoNb [80], NbTa [82] and WNb [83], elastic constants were obtained by measuring sound velocities using the pulse superposition method, and for MoTa elastic constants were obtained by measuring sound velocities using the thin rod resonance technique. High-frequency methods are generally considered to have higher accuracy than static or low frequency methods [81]. Excellent agreement between the computed rigid-ion elastic coefficients and available experimental data is observed, confirming the accuracy of the rigid-ion method for computing the elastic tensors in the present class of binary alloys. It should be noted that our electronic structure calculations are performed at 0 K, while the experimental results are obtained at 298 K; however, refractory metals/alloys are known for the stability of their mechanical properties in this temperature range [80, 81, 83, 84].
@@ 3.3 Derived mechanical properties
Polycrystalline elastic moduli were derived from the set of computed, Voigt-averaged rigid-ion elastic constants ($\bar{C}_{11}^{rigid}$, $\bar{C}_{12}^{rigid}$ and $\bar{C}_{44}^{rigid}$) for the binary alloys as a function of composition. The Voigt method [33] was used to calculate the polycrystalline elastic moduli. Fig. 2 summarizes $B$, $G$ and $E$ for all refractory binary alloys along with experimental values of $E$ [86, 87, 88, 89]. Similar to the behavior observed for the elastic constants, same-group binary alloys exhibit a linear trend as a function of composition (Figs. 2(c) and 2(d)) for all elastic moduli. While $E>B$ for all MoW compositions, $E<B$ in NbTa across all compositions. Among different-group binary alloys, the bulk modulus is observed to vary linearly with composition, while $G$ and $E$ display nonlinear behavior. Similar to $\bar{C}_{11}^{rigid}$, $\bar{C}_{12}^{rigid}$ and $\bar{C}_{44}^{rigid}$, the deviation from linearity for $G$ and $E$ is greater in case of the Ta-containing binary alloys (Figs. 2(b) and (f)) than for the Nb-containing binary alloys (Figs. 2(a) and (e)).
@@
Figure 3: Composition-dependent Pugh’s ratio ($B/G$; green) and Cauchy pressure (($C_{12}-C_{44}$); blue) for (a) MoNb; (b) MoTa; (c) MoW; (d) NbTa; (e) WNb; and (f) WTa. Horizontal lines indicate maximum Cauchy pressure or Pugh’s ratio of the constituent elements.
@@
Fig. 3 shows Pugh’s ratio ($B/G$) and the Cauchy pressure ($(C_{12}-C_{44})/2$) for the binary alloys as a function of composition. As noted above, Pugh’s ratio $B/G$ [90] and the Cauchy pressure [91, 92] are widely-used to anticipate the ductility of an alloy. A material is expected to exhibit brittle behavior if Pugh’s ratio is less than 2 [93, 94, 95], and is intrinsically ductile if the Cauchy pressure is positive [91, 92]. Similar to the behavior observed for the elastic constants and elastic moduli, MoW and NbTa show a linear trend as a function of composition for both Pugh’s ratio and the Cauchy pressure (Figs. 3(c) and 3(d)). Among different-group binary alloys, Nb-containing binary alloys displayed different behavior for Pugh’s ratio and the Cauchy pressure as a function of composition, compared to the Ta-containing binary alloys. In the case of the Nb-containing binary alloys MoNb and WNb, the $B/G$ value is always less than that of the constituent elements across the entire composition range ((Figs. 3(a) and 3(e)). However, the Cauchy pressures in MoNb and WNb have higher values than those of the constituent elements at 12.5 at.Nb composition in both binary alloys. In the case of the Ta-containing binary alloys MoTa and WTa, both $B/G$ and the Cauchy pressure values are observed to be greater than the values for their constituent elements for some compositions (Figs. 3(b) and 3(f)). In MoTa, the $B/G$ and Cauchy pressure values are greater than that of the ductile constituent element Ta up to $\approx 35$ at.and $\approx 40$ at.of Ta, respectively. In WTa, the $B/G$ and Cauchy pressure values are greater than their ductile constituent element Ta up to $\approx 37.5$ and $\approx 52.5$ at.of Ta, respectively.
3.4 Elastic anisotropy as a function of composition
@@
The finite relaxations and standard deviation values for the binary alloys summarized in Table 1 reflect different degrees of elastic anisotropy in these materials. In order to quantitate the differences among the binaries, at different compositions, we computed the $A_{G}$ and $A_{E}$ metrics defined in Eqs. (14)and (15). Table  2 compares the calculated anisotropies for the averaged relaxed-ion elastic constants. In the case of an isotropic bulk material, these values will be unity, and the deviation from unity corresponds to the degree of anisotropy in the material. The anisotropy measures $A_{G}$ and $A_{E}$ are close to one in all binary alloys, corresponding to near-isotropic behavior. These anisotropies reflect the local anisotropy at the atomic level, corresponding to different atomic bond strengths in different crystallographic planes. Although the values of $A_{G}$ and $A_{E}$ suggest that the binary alloys are near-isotropic, these measures do not provide insight into the spatial variations of the elastic moduli. Since rigid-ion elastic constants are uniform across the lattice and relaxation values are added as a correction to obtain relaxed-ion elastic constants, we therefore define a new metric, $A_{G}^{R}$. This characterizes the average change in the shear modulus value at different atomic sites across the lattice:
$A_{G}^{R}=1-\frac{\bar{G}^{\text{nuclear}}}{\bar{G}^{\text{relaxed}}}.$ (21)
@@
The values of $A_{G}^{R}$ are also listed in Table 2.
Table 2: Comparison of $A_{G}$ (Eq. (14)), $A_{E}$ (Eq. (15)) and $A_{G}^{R}$ (Eq. (21)) for all binary alloys and different compositions.
Composition $A_{G}$ $A_{E}$ $A_{G}^{R}$ Composition $A_{G}$ $A_{E}$ $A_{G}^{R}$
AB at.of A AB at.of A
MoNb 25 1.032 1.049 0.979 MoTa 25 1.007 0.987 0.939
50 1.000 1.000 0.991 50 1.000 1.000 0.977
75 0.984 0.959 0.996 75 0.996 1.015 0.995
MoW 25 1.000 1.000 0.999 NbTa 25 1.009 1.003 0.986
50 1.000 1.000 0.999 50 1.000 1.000 0.961
75 1.001 1.003 1.000 75 1.013 1.004 0.973
WNb 25 0.982 0.970 0.992 WTa 25 0.999 1.000 0.991
50 1.000 1.000 0.992 50 1.000 1.000 0.993
75 0.997 1.004 0.997 75 0.995 1.004 0.994
@@
To further analyze the extent of elastic tensor property anisotropy in the refractory binary alloys, 3D plots of the orientation-dependent $G$, $E$ and $\nu$, for three compositions each, were generated using the ELATE tool [96]. The results for MoNb and WTa are shown in Figs. 4 and 5; the results for MoTa, MoW, NbTa, and WNb can be found in the Appendix. The axes of the 3D plots in the figures correspond to the crystal orientations [1 0 0]delimited-[]100[1\,0\,0][ 1 0 0 ], [0 1 0]delimited-[]010[0\,1\,0][ 0 1 0 ], [0 0 1]delimited-[]001[0\,0\,1][ 0 0 1 ] in the $x$, $y$, and $z$ directions, respectively. The maximum and minimum values of $G$ and $\nu$ are represented as transparent blue and solid green surfaces. The extent of anisotropy is seen to be greatest for MoNb. For the Young’s modulus $E$, [1 0 0]delimited-[]100[1\,0\,0][ 1 0 0 ] is the stiffest direction and [1 1 1]delimited-[]111[1\,1\,1][ 1 1 1 ] is the most compliant direction in all compositions of the binary alloys except for $\rm{Nb_{0.25}Ta_{0.75}}$ and $\rm{W_{0.25}Ta_{0.75}}$, which have nearly perfect spherical shapes, corresponding to elastic isotropy. $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{W_{0.25}Nb_{0.75}}$ and $\rm{Nb_{0.75}Ta_{0.25}}$ deviate strongly away from spherical, indicating high elastic anisotropy. Interestingly, WTa exhibits near-isotropic behavior at all compositions as seen in Fig. 5. Similar trends are observed for $G$ and $\nu$. The patterns observed in the 3D plots are consistent with the reduced metrics $A_{E}$ and $A_{G}$ in Table 2. For $A_{E}$, all binaries showed near-isotropic behavior with deviations away from 1 of $\leq 1\%$ except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.75}Nb_{0.25}}$ and $\rm{W_{0.25}Nb_{0.75}}$, where the deviations are $>3\%$. For $A_{G}$, all binaries exhibited near-isotropic behavior (deviations of $<1\%$), except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.75}Nb_{0.25}}$, $\rm{Nb_{0.75}Ta_{0.25}}$ and $\rm{W_{0.25}Nb_{0.75}}$, which exhibit greater deviations away from unity. For $A_{G}^{R}$, all binaries showed near-isotropic behavior (deviations of $<1\%$), except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.25}Ta_{0.75}}$, $\rm{Mo_{0.5}Ta_{0.5}}$, $\rm{Nb_{0.25}Ta_{0.75}}$ and $\rm{Nb_{0.75}Ta_{0.25}}$.
@@
Figure 4: Spatial dependence of $G$ (GPa), $E$ (GPa), and $\nu$ for: (a)-(c) $\rm{Mo_{0.25}Nb_{0.75}}$;  (d)-(f) $\rm{Mo_{0.5}Nb_{0.5}}$; and (g)-(h) $\rm{Mo_{0.75}Nb_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 5: Spatial dependence of $G$ (in GPa), $E$ (in GPa), and $\nu$ for: (a)-(c) $\rm{W_{0.25}Ta_{0.75}}$;  (d)-(f) $\rm{W_{0.5}Ta_{0.5}}$;  and (g)-(h) $\rm{W_{0.75}Ta_{0.25}}$. Plots generated using the ELATE tool [96]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.

### v4

@@ 3.2 Rigid-ion elastic constants
Despite the lattice distortions in the binary alloys, the magnitudes of the relaxations lie within the range of the standard deviations and within 2% of the rigid-ion elastic constant values. These differences lie within experimental errors, which can be as high as 10% [davies1977transducer, beg1976study], depending on the experimental conditions and technique employed. Hence, the computationally faster rigid-ion elastic tensor calculations were performed for the additional compositions of A = 12.5 at.%, 37.5 at.%, 62.5 at.% and 87.5 at.% for all of the AB binary alloys. This allowed us to capture the effects of alloying on the elastic constants of BCC refractory binary alloys for an extended set of compositions.
@@
Figure 1: Composition-dependent Voigt-averaged rigid-ion elastic constants $\bar{C}_{11}^{rigid}$, $\bar{C}_{12}^{rigid}$ and $\bar{C}_{44}^{rigid}$ (solid lines, Redlich-Kister polynomial fit to computed values (filled diamonds)), compared with experimental elastic constants (stars) for (a) MoNb [hubbell1972elastic]; (b) MoTa [armstrong1970influence]; (c) MoW; (d) NbTa [fisher1981effects]; (e) WNb [frey1978elastic]; and (f) WTa [anderson1980single]. Computed error bars correspond to standard deviation with respect to Voigt averages (Eq. (2.3)).
@@
Fig. 1 compares the calculated rigid-ion elastic constants ($\bar{C}_{11}^{rigid},$ $\bar{C}_{12}^{rigid},$ and $\bar{C}_{44}^{rigid}$) for all compositions (filled diamonds) with experimental values (stars) for the binary alloys. The rigid-ion elastic constants are fitted to the Redlich-Kister polynomial given in Eq. (2.3) (solid line). For the same-group binary alloys (Figs. 1(c) and (d)), elastic constants are observed to vary linearly, similar to the empirical rule of Vegard’s law [denton1991vegard], while nonlinear behavior is observed for the different-group binary alloys. For the different-group binary alloys, the Ta-containing binary alloys (Figs. 1(b) and (f)) exhibit a greater degree of nonlinearity compared to the Nb-containing binary alloys (Figs. 1(a) and 1(e)). For $\bar{C}_{11}^{rigid}$ and $\bar{C}_{44}^{rigid}$ in the different-group binary alloys, a negative deviation from Vegard’s law is observed for the compositions with ${\rm X}>50$ at.% where X is the element with lower elastic modulus value of the constituent elements. However, for $\bar{C}_{12}^{rigid}$, the trend is reversed, i.e., a positive deviation from Vegard’s law is observed for the compositions when ${\rm X}>50$ at.% where X is the element with lower elastic modulus value of the constituent elements.
@@
Figure 2: Composition-dependent elastic moduli $B$, $G$ and $E$ computed using the Voigt method (Eqs. (10)–(12)), as a function of composition, compared with experimental $E$ (circles) for (a) MoNb [mccoy1964mechanical, campbell2008elements, WebElements-website]; (b) MoTa [mccoy1964mechanical, campbell2008elements, WebElements-website, khakurel2021machine]; (c) MoW [mccoy1964mechanical, campbell2008elements, WebElements-website]; (d) NbTa [mccoy1964mechanical, campbell2008elements, WebElements-website, khakurel2021machine]; (e) WNb [mccoy1964mechanical, campbell2008elements, WebElements-website]; and (f) WTa [mccoy1964mechanical, campbell2008elements, WebElements-website, khakurel2021machine]. Error bars correspond to the standard deviation about averaged experimental values.
@@ 3.2.1 Comparison with experiment
Experimental values at room temperature for all elastic constants were obtained from the literature as measured using high-frequency ultrasonic methods. For MoNb [hubbell1972elastic], NbTa [fisher1981effects] and WNb [frey1978elastic], elastic constants were obtained by measuring sound velocities using the pulse superposition method, and for MoTa elastic constants were obtained by measuring sound velocities using the thin rod resonance technique. High-frequency methods are generally considered to have higher accuracy than static or low frequency methods [armstrong1970influence]. Excellent agreement between the computed rigid-ion elastic coefficients and available experimental data is observed, confirming the accuracy of the rigid-ion method for computing the elastic tensors in the present class of binary alloys.
It should be noted that our electronic structure calculations are performed at 0 K, while the experimental results are obtained at 298 K; however, refractory metals/alloys are known for the stability of their mechanical properties in this temperature range [hubbell1972elastic, armstrong1970influence, frey1978elastic, anderson1980single]. Experimental data for the temperature dependence of the elastic constants $C_{11}$, $C_{12}$ and $C_{44}$ is available for the bulk elements Mo, Ta, and W [featherston1963elastic], Nb [hubbell1972elastic], and—as a function of composition—for the binary alloys MoNb [hubbell1972elastic] and WTa [anderson1980single]. The values of all elastic constants decrease monotonically as a function of temperature except for $C_{44}$ of MoNb, which increases with temperature for several compositions [hubbell1972elastic]. For the bulk elements, the differences between effective 0 K (measured 4.2 K) and 298 K $C_{11}$ and $C_{12}$ values are all $2.2\%$. For $C_{44}$, the differences between the 4.2 K and 298 K values range from 1.5% for W to 6.3% for Ta. For MoNb, the differences between measured values at 83 K and 298 K for both $C_{11}$ and $C_{12}$ are $3\%$ for all reported compositions [hubbell1972elastic]. For $C_{44}$, the differences in values between 83 K and 298 K range from $1\%$ to 13.9% (where the magnitude of $C_{44}$ is small) depending on composition. For WTa, the differences between measured values at 83 K and 298 K for both $C_{11}$ and $C_{12}$ are $2\%$ for all reported compositions [anderson1980single]. For $C_{44}$, the differences in values between 83 K and 298 K range from $1\%$ to 6.5% depending on composition. Overall, these temperature variations lie within expected experimental error bars of $\approx$ 10 % for elastic constant measurements [davies1977transducer, beg1976study].
@@ 3.3 Derived mechanical properties
Polycrystalline elastic moduli were derived from the set of computed, Voigt-averaged rigid-ion elastic constants ($\bar{C}_{11}^{rigid}$, $\bar{C}_{12}^{rigid}$ and $\bar{C}_{44}^{rigid}$) for the binary alloys as a function of composition. The Voigt method [simmons1971single] was used to calculate the polycrystalline elastic moduli. Fig. 2 summarizes $B$, $G$ and $E$ for all refractory binary alloys along with experimental values of $E$ [mccoy1964mechanical, campbell2008elements, WebElements-website, khakurel2021machine]. Similar to the behavior observed for the elastic constants, same-group binary alloys exhibit a linear trend as a function of composition (Figs. 2(c) and 2(d)) for all elastic moduli. While $E>B$ for all MoW compositions, $E<B$ in NbTa across all compositions. Among different-group binary alloys, the bulk modulus is observed to vary linearly with composition, while $G$ and $E$ display nonlinear behavior. Similar to $\bar{C}_{11}^{rigid}$, $\bar{C}_{12}^{rigid}$ and $\bar{C}_{44}^{rigid}$, the deviation from linearity for $G$ and $E$ is greater in case of the Ta-containing binary alloys (Figs. 2(b) and (f)) than for the Nb-containing binary alloys (Figs. 2(a) and (e)).
@@
Figure 3: Composition-dependent Pugh’s ratio ($B/G$; green) and Cauchy pressure $($$C_{12}-C_{44}$); blue) for (a) MoNb; (b) MoTa; (c) MoW; (d) NbTa; (e) WNb; and (f) WTa. Horizontal lines indicate maximum Cauchy pressure or Pugh’s ratio of the constituent elements.
@@
Fig. 3 shows Pugh’s ratio ($B/G$) and the Cauchy pressure ($(C_{12}-C_{44})/2$) for the binary alloys as a function of composition. As noted above, Pugh’s ratio $B/G$ [pugh1954xcii] and the Cauchy pressure [pettifor1992theoretical, pettifor1991bonding] are widely-used to anticipate the ductility of an alloy. A material is expected to exhibit brittle behavior if Pugh’s ratio is less than 2 [gilman2003electronic, niu2012extra, thompson2018predicting], and is intrinsically ductile if the Cauchy pressure is positive [pettifor1992theoretical, pettifor1991bonding]. Similar to the behavior observed for the elastic constants and elastic moduli, MoW and NbTa show a linear trend as a function of composition for both Pugh’s ratio and the Cauchy pressure (Figs. 3(c) and 3(d)). Among different-group binary alloys, Nb-containing binary alloys displayed different behavior for Pugh’s ratio and the Cauchy pressure as a function of composition, compared to the Ta-containing binary alloys. In the case of the Nb-containing binary alloys MoNb and WNb, the $B/G$ value is always less than that of the constituent elements across the entire composition range ((Figs. 3(a) and 3(e)). However, the Cauchy pressures in MoNb and WNb have higher values than those of the constituent elements at 12.5 at.% Nb composition in both binary alloys. In the case of the Ta-containing binary alloys MoTa and WTa, both $B/G$ and the Cauchy pressure values are observed to be greater than the values for their constituent elements for some compositions (Figs. 3(b) and 3(f)). In MoTa, the $B/G$ and Cauchy pressure values are greater than that of the ductile constituent element Ta up to $\approx 35$ at.% and $\approx 40$ at.% of Ta, respectively. In WTa, the $B/G$ and Cauchy pressure values are greater than their ductile constituent element Ta up to $\approx 37.5$ and $\approx 52.5$ at.% of Ta, respectively.
3.4 Elastic anisotropy and elastic heterogeneity as a function of composition
@@
The finite relaxations and standard deviation values in the elastic constants for the binary alloys summarized in Table 1 reflect different degrees of spatial variations (elastic heterogeneity) in these materials. In order to quantitate the differences in elastic anisotropies among the binaries, at different compositions, we computed the $A_{G}$ and $A_{E}$ metrics defined in Eqs. (14) and (15). Table  2 compares the calculated anisotropies for the relaxed-ion elastic constants. In the case of an isotropic bulk material, these values will be unity, and the deviation from unity corresponds to the degree of anisotropy in the material. The anisotropy measures $A_{G}$ and $A_{E}$ are close to one in all binary alloys, corresponding to near-isotropic behavior. These anisotropies reflect the local anisotropy at the atomic level, corresponding to different atomic bond strengths in different crystallographic planes. Although the values of $A_{G}$ and $A_{E}$ suggest that the binary alloys are near-isotropic, these measures do not provide insight into the spatial variations of the elastic moduli. Since rigid-ion elastic constants are uniform across the lattice and relaxation values are added as a correction to obtain relaxed-ion elastic constants, we therefore define two new metrics, $R_{G}$ and $R_{E}$, expressed in terms of the respective ratios of the nuclear and relaxed contributions. An elastic heterogeneity constant with a value of unity indicates the absence of spatial variation in elastic properties. These characterize the average changes in the shear modulus and Young’s modulus values at different atomic sites across the lattice:
${R_{G}}=1-\frac{G_{\text{nuclear}}}{G_{\text{relaxed}}}$ (38)
and
$R_{E}=1-\frac{E_{\text{nuclear}}}{E_{\text{relaxed}}},$ (39)
@@
where $G_{\text{nuclear}}$ and $G_{\text{relaxed}}$ are defined in Eqs. (2.4) and (20) and $E_{\text{nuclear}}$ and $E_{\text{relaxed}}$ are defined in Eqs. (25) and (27). The values of $R_{G}$ and $R_{E}$ are also listed in Table 2.
Table 2: Comparison of $A_{G}$ (Eq. (14)) with $R_{G}$ (Eq. (38)) and $A_{E}$ (Eq. (15)) with $R_{E}$ (Eq. (39)) and for all binary alloys and different compositions.
Composition $A_{G}$ $R_{G}$ $A_{E}$ $R_{E}$ Composition $A_{G}$ $R_{G}$ $A_{E}$ $R_{E}$
AB at.% of A AB at.% of A
MoNb 25 1.032 0.979 1.049 0.981 MoTa 25 1.007 0.939 0.987 0.945
50 1.000 0.991 1.000 0.991 50 1.000 0.977 1.000 0.979
75 0.984 0.996 0.959 0.996 75 0.996 0.995 1.015 0.995
MoW 25 1.000 0.999 1.000 0.999 NbTa 25 1.009 0.986 1.003 0.988
50 1.000 0.999 1.000 0.999 50 1.000 0.961 1.000 0.966
75 1.001 1.000 1.003 0.999 75 1.013 0.973 1.004 0.976
WNb 25 0.982 0.992 0.970 0.992 WTa 25 0.999 0.975 1.000 0.976
50 1.000 0.992 1.000 0.991 50 1.000 0.982 1.000 0.983
75 0.997 0.997 1.004 0.997 75 0.995 0.994 1.004 0.997
@@
To further analyze the extent of elastic tensor property anisotropy in the refractory binary alloys, 3D plots of the orientation-dependent $G$, $E$ and $\nu$, for three compositions each, were generated using the ELATE tool [gaillac2016elate]. The results for MoNb and WTa are shown in Figs. 4 and 5; the results for MoTa, MoW, NbTa, and WNb can be found in the Appendix. The axes of the 3D plots in the figures correspond to the crystal orientations [1 0 0][1\,0\,0], [0 1 0][0\,1\,0], [0 0 1][0\,0\,1] in the $x$, $y$, and $z$ directions, respectively. The maximum and minimum values of $G$ and $\nu$ are represented as transparent blue and solid green surfaces. The extent of anisotropy is seen to be greatest for MoNb. For the Young’s modulus $E$, [1 0 0][1\,0\,0] is the stiffest direction and [1 1 1][1\,1\,1] is the most compliant direction in all compositions of the binary alloys except for $\rm{Nb_{0.25}Ta_{0.75}}$ and $\rm{W_{0.25}Ta_{0.75}}$, which have nearly perfect spherical shapes, corresponding to elastic isotropy. $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{W_{0.25}Nb_{0.75}}$ and $\rm{Nb_{0.75}Ta_{0.25}}$ deviate strongly away from spherical, indicating high elastic anisotropy. Interestingly, WTa exhibits near-isotropic behavior at all compositions as seen in Fig. 5. Similar trends are observed for $G$ and $\nu$. The patterns observed in the 3D plots are consistent with the reduced metrics $A_{E}$ and $A_{G}$ in Table 2. For $A_{E}$, all binaries showed near-isotropic behavior with deviations away from 1 of $\leq 1\%$ except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.75}Nb_{0.25}}$ and $\rm{W_{0.25}Nb_{0.75}}$, where the deviations are $3\%$. For $A_{G}$, all binaries exhibited near-isotropic behavior (deviations of $1\%$), except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.75}Nb_{0.25}}$, $\rm{Nb_{0.75}Ta_{0.25}}$ and $\rm{W_{0.25}Nb_{0.75}}$, which exhibit greater deviations away from unity. For both $R_{G}$ and $R_{E}$, all binaries showed near-isotropic behavior (deviations of $1\%$), except for $\rm{Mo_{0.25}Nb_{0.75}}$, $\rm{Mo_{0.25}Ta_{0.75}}$, $\rm{Mo_{0.5}Ta_{0.5}}$, $\rm{Nb_{0.25}Ta_{0.75}}$, $\rm{Nb_{0.5}Ta_{0.5}}$, $\rm{Nb_{0.75}Ta_{0.25}}$, $\rm{W_{0.25}Ta_{0.75}}$ and $\rm{W_{0.5}Ta_{0.5}}$.
@@
Figure 4: Anisotropy plots of $G$ (GPa), $E$ (GPa), and $\nu$ for: (a)-(c) $\rm{Mo_{0.25}Nb_{0.75}}$;  (d)-(f) $\rm{Mo_{0.5}Nb_{0.5}}$; and (g)-(h) $\rm{Mo_{0.75}Nb_{0.25}}$. Plots generated using the ELATE tool [gaillac2016elate]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.
@@
Figure 5: Anisotropy plots of $G$ (in GPa), $E$ (in GPa), and $\nu$ for: (a)-(c) $\rm{W_{0.25}Ta_{0.75}}$;  (d)-(f) $\rm{W_{0.5}Ta_{0.5}}$;  and (g)-(h) $\rm{W_{0.75}Ta_{0.25}}$. Plots generated using the ELATE tool [gaillac2016elate]. The maximum and minimum values of $G$, $E$ and $\nu$ are represented as transparent blue and solid green surfaces respectively.